OpenStudy (anonymous):

How does the graph of f (x) = ( x -1)^2 relate to the basic function f (x)=x^2 A) shift 1 unit to the left B) shift 1 unit to the right C)shift 1 until up D) shift 1 until down

OpenStudy (campbell_st):

have you graphed the curves..?

OpenStudy (campbell_st):

can you identify the zeros of each curve..?

OpenStudy (ybarrap):

$$ y=(x-h)^2 $$ Will shift \(y=x^2\) to the RIGHT by h units $$ y=(x+h)^2 $$ Will shift \(y=x^2\) to the LEFT by h units Does this help?